Prev Next

IT Weekly Meeting Minutes 2020-05-12

Attendees: Stephen, David, LeRoy, Matthieu

Discussion

  • new fiber run updates / via pepsi cable chain
    • LR digging into unterminated fibers we have at the moment (existing run)
    • still need to verify the actual run length
  • synology disks SMART failing
    • issue was fixed last week when SH did the replacement
    • reboot today, wait some time if the emails stop
  • LUCI upgrade plan / parts
    • with LR, confident we know what's inside the machine
    • one/two kits ($700/per unit - raid controllers + cable + SSDs)
      • how old are those machines? PCIe.
      • keep the OS the same, CentOS-7? nope. 6 or 5. losely decided to use 7 with a conversation with IS.
      • started because we got S.M.A.R.T warnings on the machine.
      • currently soft-raid array
      • cost new vs. uprgade.
        • we have some DELL's lying around we could repurpose - they're 1U. We need 2x full height slots for the CCD controllers. Check if the spare DELLs can meet that.
      • how much a new server? $3000 a piece, they're 7 years old.
      • how about a new server for luci, spare coming from a general pool of spares? we'd keep the "old" lucix as spare.
      • DE will take another crack at it
  • new KVM / java
    • 3 clients
      • Active Client (.NET)
      • VKC vitual kvm client (java, but need the oracle, not open-java)
      • HTML/javascript client (does not support all the features)
    • multiple connection? there is a timeout option. stop moving the mouse, allows someone else to get in.
    • should try it out.
    • multiple viewers still possible
    • local seat needs a keyboard and screen
    • USB connection (would have worked from windows? didn't seem to)
    • fine for Tucson but maybe not for the mountain
  • ROWG updates
    • new screens
      • 32" arrive being setup
        • setup brand new machine "507-2"
        • old 507-2 is 507-old (32 bits)
      • running out of host hardware / space for the 2+1 new ones.
      • 2+1 other ones being ordered - use existing rm507
      • Petr's office for the big screen tests. Leave the key in the printer room or something
    • 507 distancing options
  • FOSWiki landing page recovery
    • how to avoid the issue in the future
    • bigger question of incremental backup
    • DE will create a sheet of what is google backed up
  • CR power-on Mon 18
    • PEPSI
    • archive
    • luci
    • sx-AO
    • iLocater trunking
    • TCS (get rid of tcs-test for telemetry)
      • mods/lbc will stay down for now
    • AOC - work on AGW4, go up
    • save us some rush if there's a restart
  • weather station status
    • Ramon fixed it with LR
  • server / OS inventory
  • CR temperature sensors
    • they came back from the moutain, send back again
  • mac-mini
    • in DE office
  • JW printer
    • DE will send a proposal, USB+Ethernet
  • AOB

Topic revision: r3 - 19 May 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback